OpenStudy (ambermarie151):

Complete the square to solve 4x^2+24x=4

OpenStudy (ambermarie151):

@will.h

OpenStudy (will.h):

\[4x^2 + 24x = 4\] You can get all the terms on one side and set the equation equal to 0 like the following \[4x^2 + 24x - 4 = 0\] Now you can factor or use the famous quadratic formula let's use famous quadrtaic formula \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] we have b = 24 and a = 4 and c = -4 you can take it from here

OpenStudy (jango_in_dtown):

hiii @ambermarie151

OpenStudy (ambermarie151):

So when I plug then in it will be 24-4=20 then divide by 2 which equals 10 @will.h

OpenStudy (jango_in_dtown):

first of all divide both sides by 4

OpenStudy (jango_in_dtown):

then you get x^2+6x=1

OpenStudy (jango_in_dtown):

then add both sides 9

OpenStudy (jango_in_dtown):

then you get (x+3)^2=10

OpenStudy (ambermarie151):

I am so confused right now lol

OpenStudy (jango_in_dtown):

i.e. x=-3+ sqrt(10) and -3-sqrt(10)

OpenStudy (will.h):

do you see the \[\pm\]? it means that we will have 2 answers the 1st when you apply the rule when it is + and the 2nd when it is - i will help you when it is + and you do the exact same thing for when it is negative just change the sign in between the sqrt and -b so here's the 1st \[\frac{ -24 + \sqrt{24^2 - 4(4)(-4)} }{ 2*4 }\] So the 1st solution is 0.162277 The 2nd (when it is negative) = -6.1623

OpenStudy (will.h):

Hope that helps.. i have a test i have to go.. Bye and goodl luck

OpenStudy (jango_in_dtown):

@Will.H thats not the method of completing ssquares

OpenStudy (will.h):

@ambermarie151 here's the proof in the attachment and jango There are a thousand ways to solve this and pretty sure Famous Quadratic formula is the most famous way..

OpenStudy (jango_in_dtown):

Thats Sridhar Acharya's method what you used, but thats not the method of completing squares. The question said to solve by completing squares @Will.H

OpenStudy (will.h):

yeah i hadn't notice sorry but if it is multiply choices then you haad the answer if not then Jango will help he seems to know about the completing the square don't you jango?

OpenStudy (jango_in_dtown):

yes the method I did is called the method of completing squares, when you have to express the term in form of some squares and then give the final answer